OpenStudy (anonymous):

A golf ball is hit from the ground with an initial speed of 192 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground? (1 point) 39 secs 19 secs 12 secs 6 secs

OpenStudy (anonymous):

use the motion equations of constant acceleration s = ut + 0.5at^2 displacement s = 0, a = -32ft/sec^2 and u = 192 ft/s 0 = 192t - 16t^2 solving for t will give the answer

OpenStudy (anonymous):

can you do this?

OpenStudy (anonymous):

I can try :p

OpenStudy (anonymous):

on root wil be zero which we can ignore

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im still confused how you solve for t

OpenStudy (anonymous):

ok first factor 0 = t(192 - 16t) either t = 0 or 192 - 16t = 0 16t = 192 t = 192 / 16 = 12 12 seconds

OpenStudy (anonymous):

oh ok thanks!

OpenStudy (anonymous):

your welcome